# Age Group Sensitivity Analysis of Epidemic Models: Investigating the Impact of Contact Matrix Structure

### v1

@@ 2.4 Partial Rank Correlation Coefficient (PRCC)
The Partial Rank Correlation Coefficient (PRCC) is a widely-used tool for sensitivity analysis based on Latin Hypercube Sampling (LHS). Pearson’s correlation coefficient, partial correlation coefficient, and standardized regression coefficients assume linear relationships between variables. In contrast, the rank correlation variants test for non-linear but monotonic relationships. This method works by replacing actual values with rank numbers, thereby mitigating the effect of non-linearity in the relationship. Examples of rank correlation methods include SRCC (Spearman Rank Correlation Coefficient), PRCC, and SRRC (Standardized Rank Regression Coefficient) [3, 6, 11].
The initial step in calculating the PRCC values involves replacing each LHS value in Table 1 with its rank value. A rank value is an integer if the corresponding data point is unique. If multiple identical values exist in the sample, they are assigned the average of their respective ranks. Ranking is executed for every parameter and the simulation output values, as illustrated in Table 2. We use $r_{i}$ ($i=1,2,\dots,K$) to represent the input parameters with ranks and $d$ for the target variable. Subsequently, the procedure involves fitting $2K$ regression models in two rounds: using $d$ as a target in the first round and then fitting models with each $r_{i}$ as the output variable. In the second round of linear regression, the rank parameter $r_{k}$ is estimated using the other rank parameters [5]. In neither round of the linear regression process is the parameter $r_{k}$ used as an input variable.
@@ 2.5 Inference on PRCCs
Marino et al. [5] describe a technique to determine the significance of tests that measure whether a pairwise PRCC value $P_{k}$ is significantly different from zero and whether two pairwise PRCC values differ significantly from each other. As per the authors, if the sample size is $N$ and there are $K$ parameters whose effects are considered when calculating $P_{k}$, each pairwise PRCC ($P_{k}$) will yield a value $\mathcal{T}_{k}$ according to the following statistic:
@@
The 136 sensitivity measures for $P_{k}$ and their corresponding $p_{k}$, are arranged in a square symmetric matrix to be used in the aggregation approach. A lower triangular of a matrix like this can be seen on Fig. 3. For each $P_{k}$ element, we determine the corresponding $(i,j)$ age group pair to which the sensitivity value belongs. This value is then assigned to the $P_{i,j}$ and $P_{j,i}$ elements of the $P$ matrix. We repeat this process for the $p_{k}$ values as well, assigning each to the corresponding $p_{i,j}$ and $p_{j,i}$ elements of the $p$ matrix. Consequently, we define the matrices $P$ and $p$ as follows:
@@ 2.6 PRCC aggregation approach
Now, we introduce the methodology for aggregating sensitivity values across age group pairs. For a given $(i,j)$ contact pair, the sensitivity value $P_{i,j}$ is associated with the corresponding $p_{i,j}$ value, which represents the uncertainty of the sensitivity measure. Thus, $1-p_{i,j}$ quantifies the reliability of $P_{i,j}$. Now, we construct a probability distribution for age group $i$ using the values of $1-p_{i,j}$.
@@ 3 Demonstrations
To illustrate the utility and versatility of the AGSA method, we applied it to two age-structured epidemic model with different complexity. In this section, we briefly introduce these models and present the aggregated sensitivity values in case of various target parameters. For further details of the models, we refer the reader to the Supplementary Material and the original paper [32, 36]. Using the LHS method, the parameter intervals were divided into 10,000 parts, representing the sample size. Simulations were conducted until the number of infected individuals dropped below one. During the simulations, we considered two scenarios: a mild epidemic with ${\mathcal{R}}_{0}=1.2$ and a severe epidemic with ${\mathcal{R}}_{0}=2.5$.
3.1 SEIR model for influenza epidemic by Pitman et al.
The first model is an SEIRS (Susceptible-Exposed-Infectious-Recovered-Susceptible) model designed to simulate the spread of influenza, incorporating age-specific mixing patterns in England and Wales [36]. The model employs a Who Acquires Infection From Whom (WAIFW) matrix, derived from the POLYMOD study [13], which categorizes interactions among 15 age groups while distinguishing between physical and non-physical contacts. This model emphasizes the critical role of pediatric vaccination in mitigating the burden of influenza, demonstrating how herd immunity can effectively protect vulnerable populations. In our simulations, we used the parameters of the original model (along with the population data of England and Wales), with the following modification: we assume no waining of immunity, thus our variant is an SEIR type model and we assume no vaccination in these simulations.
@@
In this section we demonstrate the output of the contact matrix sensitivity analysis and the output of the aggregation on Fig. 3. The left panels of this figure show the pairwise PRCC and $p$ values (according to Eq. 2.6 and 2.7) that are the output of the contact matrix sensitivity analysis. Upon aggregation, the right panel highlights age groups with the highest aggregated sensitivity values, indicating their significant influence on the model’s target variables. We observe that, in the case of a mild epidemic, the system is most sensitive to the contacts of individuals aged 35–39. However, in the case of a severe epidemic, the contacts of nine age groups (spanning between 15-59) are significant. This insight demonstrates that, for a given target variable, the sensitivity of different age groups can vary significantly depending on the severity of the epidemic.
Figure 3: Demonstrations of the AGSA framework in the Influenza model by Pitman et al. Target parameter: number of infected individuals at the peak of the epidemic. Top row: mild influenza outbreak with $\mathcal{R}_{0}=1.2$, bottom row: severe outbreak with $\mathcal{R}_{0}=2.5$. In both cases, the target variable is the number of infected individuals at the peak of the epidemic. Left panels: pairwise PRCC values and associated p-values for the sensitivity analysis of contact matrix elements. Notice that both the diagonal of the PRCC matrix and the p-value matrix (Eq. 2.6 and 2.7) are shown on these panels. Right panels: aggregated PRCC values by age group.
@@
The model [32] we use for the demonstrations is a detailed age-structured COVID-19 model developed for Hungary, incorporating Erlang-like distribution of latent and infectious period, various types of hospitalisation and death compartments. The contact data for Hungary is from Prem et al.[26] and the population vector is obtained from the Hungary Central Statistical Office (KSH). The model incorporates 16 age groups and 15 specific disease related classes per age group.
@@
This model is sufficiently complex to define several target functions. We analyze the sensitivity of contacts with respect to the basic reproduction number, ICU peak and cumulative fatalities under both mild and severe epidemic scenarios.
A brief note on how the basic reproduction number is used as a target variable: for a given model and set of parameters, the reproduction number is calculated as described in Section 2.3 based on Diekmann et. al [35]. The desired $\mathcal{R}_{0}$ values for each simulation run are then scaled using the transmission parameter $\beta$, and simulations are conducted with these settings. During the LHS procedure, sampling the contact matrix naturally leads to variations in the set $\mathcal{R}_{0}$ value. Therefore, $\mathcal{R}_{0}$ can also be treated as a target variable within the method. The summarized results are presented in Fig. 4.
@@
Figure 4: Aggregated sensitivity values ($P$) and corresponding confidence intervals (CI) across age groups for different target variables in mild ($\mathcal{R}_{0}=1.2$, top row) and severe ($\mathcal{R}_{0}=2.5$, bottom row) epidemic scenarios. The target variables include the basic reproduction number ($\mathcal{R}_{0}$), ICU peak occupancy, and cumulative fatalities. The legend on the bottom, right figure applies for all.
@@
In case of the mild epidemic scenario ($\mathcal{R}_{0}=1.2$) the highest sensitivity values are associated with the 30–44 age groups whet the target is the reproduction number and the ICU peak. A significant sensitivity is also observed in the 15–19 age group. This highlights the critical role of high school-aged individuals in driving transmission dynamics in mild outbreaks, due to their high contact rates in school and social settings. Interestingly, the number of cumulative fatalities during the whole epidemic is most sensitive for age group 40-44, not for the elderly age group (75+).
@@
In the severe epidemic scenario ($\mathcal{R}_{0}=2.5$): The 30–44 age group remains the most influential across all target variables, showing consistently high sensitivity values. This age group likely plays a pivotal role due to its high contact rates across various settings, such as workplaces and social environments. Unlike the mild scenario, the sensitivity of the 15–19 age group significantly diminishes, suggesting that during a severe epidemic, reducing contacts within younger populations has a much smaller impact on epidemic outcomes. Not surprisingly, cumulative fatalities are strongly influenced by contacts in the 75+ age group in case of the severe epidemic scenario.
@@
The contrast between mild and severe epidemic scenarios reveals that in a mild outbreak, reducing contacts in both younger age groups (particularly 15–19) and middle-aged adults (30–44) can effectively mitigate transmission and reduce key outcome measures, including ICU occupancy and cumulative fatalities. In a severe outbreak, the system becomes less responsive to contact reductions in younger age groups. The effectiveness of interventions is primarily driven by targeting middle-aged adults (30–44 years), reflecting their significant role in maintaining transmission chains in high-transmissibility scenarios.
@@
These findings suggest that age-specific intervention strategies should be adapted based on the severity of the outbreak. The figures underscore the importance of tailoring intervention strategies to the transmissibility of the disease and highlights the potential of targeted contact reductions to minimize hospital burden and fatalities effectively.
@@ 4 Discussion
In this study we presented the Age Group Sensitivity Analysis method, a novel framework designed to quantify the sensitivity of epidemic outcomes to variations in social contact structures across age groups. By integrating age-structured epidemic models with Latin Hypercube Sampling and the Partial Rank Correlation Coefficient approach, the AGSA method enables a systematic and robust evaluation of how changes in contact patterns affect key epidemic outcomes. The proposed sensitivity aggregation technique provides a nuanced view by highlighting age groups that contribute the most to uncertainty in model predictions, which can be particularly valuable for targeted intervention strategies. The findings we presented in Sec. Demonstrations, emphasize the importance of tailoring intervention strategies based on the severity of the epidemic.
A critical implication of this research lies in its potential application to improve data collection efforts. Given the challenges associated with accurately estimating social contact matrices—especially due to varying survey responses and accessibility across age groups—the AGSA method provides a data-driven basis for focusing empirical studies on age groups that introduce the greatest epistemic uncertainty. By concentrating on these influential groups, it becomes possible to refine contact structure estimates, thus enhance model predictions.
Furthermore, the framework’s ability to incorporate various target variables—such as peak hospitalization rates, ICU occupancy, and cumulative fatalities—demonstrates its versatility and practical utility for public health decision-making. Policymakers can leverage the insights provided by AGSA to prioritize interventions that align with specific epidemic control objectives, whether the focus is on reducing transmission rates, minimizing hospital burden, or lowering mortality.
In conclusion, the AGSA framework offers a comprehensive and adaptable approach to understanding the role of age-specific contact structures in disease transmission dynamics. By identifying the most influential age groups across different epidemic scenarios, this method can inform more precise and effective public health responses. Future research could extend this framework by incorporating other demographic factors, such as socioeconomic status or health comorbidities, further enhancing its applicability to real-world epidemic management.
5 Ethics approval and consent to participate
Not applicable.
6 Consent for publication
Not applicable.
@@ 7 Data Availability
Code and sensitivity related data used for this analysis are available on Github: https://github.com/zsvizi/sensitivity-contact-epidemics
@@ 8 Acknowledgements
This research was partially suppported under the project TKP2021-NVA-09, provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.

### v2

@@ 2.4 Partial Rank Correlation Coefficient (PRCC)
The Partial Rank Correlation Coefficient (PRCC) is a widely used tool for sensitivity analysis based on Latin Hypercube Sampling (LHS). Pearson’s correlation coefficient, partial correlation coefficient, and standardized regression coefficients assume linear relationships between variables. In contrast, the rank correlation variants test for non-linear but monotonic relationships. This method replaces actual values with rank numbers, thereby mitigating the effect of non-linearity in the relationship. Examples of rank correlation methods include SRCC (Spearman Rank Correlation Coefficient), PRCC, and SRRC (Standardized Rank Regression Coefficient) [3, 6].
The initial step in calculating the PRCC values involves replacing each LHS value in Table 1 with its rank value. A rank value is an integer if the corresponding data point is unique. If multiple identical values exist in the sample, they are assigned the average of their respective ranks. The ranking is executed for every parameter and the simulation output values, as illustrated in Table 2. We use $r_{i}$ ($i=1,2,\dots,K$) to represent the target variable’s input parameters with ranks and $d$. Subsequently, the procedure involves fitting $2K$ regression models in two rounds: using $d$ as a target in the first round and then fitting models with each $r_{i}$ as the output variable. In the second round of linear regression, the rank parameter $r_{k}$ is estimated using the other rank parameters [5]. In neither round of the linear regression process is the parameter $r_{k}$ used as an input variable.
@@ 2.5 Inference on PRCCs
Marino et al. [5] describe a technique to determine the significance of tests that measure whether a pairwise PRCC value $P_{k}$ significantly differs from zero and whether two pairwise PRCC values differ significantly. As per the authors, if the sample size is $N$ and there are $K$ parameters whose effects are considered when calculating $P_{k}$, each pairwise PRCC ($P_{k}$) will yield a value $\mathcal{T}_{k}$ according to the following statistic:
@@
The 136 sensitivity measures for $P_{k}$ and their corresponding $p_{k}$, are arranged in a square symmetric matrix to be used in the aggregation approach. A lower triangular of a matrix like this can be seen in Fig. 3. For each $P_{k}$ element, we determine the corresponding $(i,j)$ age group pair to which the sensitivity value belongs. This value is then assigned to the $P_{i,j}$ and $P_{j,i}$ elements of the $P$ matrix. We repeat this process for the $p_{k}$ values as well, assigning each to the corresponding $p_{i,j}$ and $p_{j,i}$ elements of the $p$ matrix. Consequently, we define the matrices $P$ and $p$ as follows:
@@ 2.6 PRCC aggregation approach
We introduce the methodology for aggregating sensitivity values across age group pairs. For a given $(i,j)$ contact pair, the sensitivity value $P_{i,j}$ is associated with the corresponding $p_{i,j}$ value, which represents the uncertainty of the sensitivity measure. Thus, $1-p_{i,j}$ quantifies the reliability of $P_{i,j}$. Now, we construct a probability distribution for age group $i$ using the values of $1-p_{i,j}$.
@@ 3 Demonstrations
We applied the AGSA method to two age-structured epidemic models of different complexity to illustrate its utility and versatility. This section briefly introduces these models and presents the aggregated sensitivity values for various target parameters. For further details of the models, we refer the reader to the original papers [33, 37]. Using the LHS method, the parameter intervals were divided into 10,000 parts, representing the sample size. Simulations were conducted until the number of infected individuals dropped below one. During the simulations, we considered two scenarios: a mild epidemic with ${\mathcal{R}}_{0}=1.2$ and a severe outbreak with ${\mathcal{R}}_{0}=2.5$.
3.1 SEIR model for an influenza epidemic by Pitman et al.
The first model is an SEIRS (Susceptible-Exposed-Infectious-Recovered-Susceptible) model designed to simulate the spread of influenza, incorporating age-specific mixing patterns in England and Wales [37]. The model employs a Who Acquires Infection From Whom (WAIFW) matrix, derived from the POLYMOD study [13], which categorizes interactions among 15 age groups while distinguishing between physical and non-physical contacts. This model emphasizes the critical role of pediatric vaccination in mitigating the influenza burden and demonstrates how herd immunity can effectively protect vulnerable populations. In our simulations, we used the parameters of the original model (along with the population data of England and Wales), with the following modification: we assume no waining of immunity, thus our variant is an SEIR type model and we assume no vaccination in these simulations.
@@
In this section, we demonstrate the output of the contact matrix sensitivity analysis and the output of the aggregation in Fig. 3. The left panels of this figure show the pairwise PRCC and $p$ values (according to Eq. 2.6 and 2.7) that are the output of the contact matrix sensitivity analysis. Upon aggregation, the right panel highlights age groups with the highest aggregated sensitivity values, indicating their significant influence on the model’s target variables. In the case of a mild epidemic, the system is most sensitive to contacts of individuals aged 35–39. However, in the case of a severe epidemic, the contacts of nine age groups (15 to 59) are significant. This insight demonstrates that, for a given target variable, the sensitivity of different age groups can vary significantly depending on the epidemic’s severity.
Figure 3: Demonstrations of the AGSA framework in the Influenza model by Pitman et al. Target parameter: number of infected individuals at the epidemic’s peak. Top row: mild influenza outbreak with $\mathcal{R}_{0}=1.2$, bottom row: severe outbreak with $\mathcal{R}_{0}=2.5$. In both cases, the target variable is the number of infected individuals at the epidemic’s peak. Left panels: pairwise PRCC values and associated p-values for the sensitivity analysis of contact matrix elements. Notice that both the diagonal of the PRCC matrix and the p-value matrix (Eq. 2.6 and 2.7) are shown on these panels. Right panels: aggregated PRCC values by age group.
@@
The model [33] we use for the demonstrations is a detailed age-structured COVID-19 model developed for Hungary, incorporating Erlang-like distribution of latent and infectious period, various types of hospitalization, and death compartments. The contact data for Hungary is from Prem et al.[27], and the population vector is obtained from the Hungary Central Statistical Office (KSH). The model incorporates 16 age groups and 15 specific disease-related classes per age group.
@@
This model is sufficiently complex to define several target functions. We analyze the sensitivity of contacts concerning the basic reproduction number, ICU peak, and cumulative fatalities under both mild and severe epidemic scenarios. The summarized results are presented in Fig. 4.
@@
Figure 4: Aggregated sensitivity values ($P$) and corresponding confidence intervals (CI) across age groups for different target variables in mild ($\mathcal{R}_{0}=1.2$, top row) and severe ($\mathcal{R}_{0}=2.5$, bottom row) epidemic scenarios. The target variables include the basic reproduction number ($\mathcal{R}_{0}$), ICU peak occupancy, and cumulative fatalities. The legend on the bottom right figure applies to all.
@@
In the case of the mild epidemic scenario ($\mathcal{R}_{0}=1.2$), the highest sensitivity values are associated with the 30–44 age groups where the target is the reproduction number and the ICU peak. A significant sensitivity is also observed in the 15–19 age group. This highlights the critical role of high school-aged individuals in driving transmission dynamics in mild outbreaks, as they have high contact rates in school and social settings. Interestingly, the number of cumulative fatalities during the whole epidemic is most sensitive for the age group 40-44, not for the elderly age group (75+).
@@
In the severe epidemic scenario ($\mathcal{R}_{0}=2.5$): The 30–44 age group remains the most influential across all target variables, showing consistently high sensitivity values. This age group likely plays a pivotal role due to its high contact rates across various settings, such as workplaces and social environments. Unlike the mild scenario, the sensitivity of the 15–19 age group significantly diminishes, suggesting that reducing contacts within younger populations has a much smaller impact on epidemic outcomes during a severe epidemic. Not surprisingly, cumulative fatalities are strongly influenced by contacts in the 75+ age group in case of a severe epidemic scenario.
@@
The contrast between mild and severe epidemic scenarios reveals that in a mild outbreak, reducing contacts in both younger age groups (mainly 15–19) and middle-aged adults (30–44) can effectively mitigate transmission and reduce key outcome measures, including ICU occupancy and cumulative fatalities. In a severe outbreak, the system becomes less responsive to contact reductions in younger age groups. The effectiveness of interventions is primarily driven by targeting middle-aged adults (30–44 years), reflecting their significant role in maintaining transmission chains in high-transmissibility scenarios.
@@
These findings suggest that age-specific intervention strategies should be adapted based on the outbreak’s severity. The figures underscore the importance of tailoring intervention strategies to the disease’s transmissibility and highlight the potential of targeted contact reductions to effectively minimize hospital burden and fatalities.
@@ 4 Discussion
In this study, we presented the Age Group Sensitivity Analysis method, a novel framework designed to quantify the sensitivity of epidemic outcomes to variations in social contact structures across age groups. The proposed sensitivity aggregation technique provides a nuanced view by highlighting age groups that contribute the most to uncertainty in model predictions, which can be particularly valuable for targeted intervention strategies. Our findings in Sec. Demonstrations emphasize the importance of tailoring intervention strategies to the epidemic’s severity.
A critical implication of this research lies in its potential application to improve data collection efforts. Given the challenges associated with accurately estimating social contact matrices—mainly due to varying survey responses and accessibility across age groups—the AGSA method provides a data-driven basis for focusing empirical studies on age groups that introduce the greatest epistemic uncertainty. Concentrating on these influential groups makes it possible to refine contact structure estimates, thus enhancing model predictions.
Furthermore, the framework’s ability to incorporate various target variables — peak hospitalization, ICU occupancy, and cumulative fatalities — demonstrates its versatility and practical utility for public health decision-making. Policymakers can leverage the insights provided by AGSA to prioritize interventions that align with specific epidemic control objectives, whether the focus is reducing transmission rates, minimizing hospital burden, or lowering mortality.
The AGSA framework offers a comprehensive and adaptable approach to understand the role of age-specific contact structures in disease transmission dynamics. This method can inform more precise and effective public health responses by identifying the most influential age groups across different epidemic scenarios. Future research could extend this framework by incorporating other demographic factors, such as socioeconomic status or health comorbidities, further enhancing its applicability to real-world epidemic management.
5 Ethics Statement
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT-4o to enhance clarity, grammar, and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ 6 Data Availability
Code and sensitivity-related data used for this analysis are available on Github: https://github.com/zsvizi/sensitivity-contact-epidemics
@@ 7 Acknowledgements
This research was partially supported by the project TKP2021-NVA-09, provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
Appendix A The Governing Equations of the SEIR Epidemic Model by Pitman et al.
For age group $i\in\{0,\dots 14\}$ the governing equations of the SEIR disease model adopted from [37] take the form of
$\frac{dS_{i}}{dt}$ $=-\beta_{0}S_{i}I_{i},$ (A.1)
$\frac{dE_{i}}{dt}$ $=\beta_{0}S_{i}I_{i}-\sigma E_{i},$
$\frac{dI_{i}}{dt}$ $=\sigma E_{i}-\gamma I_{i},$
$\frac{dR_{i}}{dt}$ $=\gamma I_{i}.$
Variable Description
$S_{i}(t)$ Number of susceptible individuals in age group $i$
$E_{i}(t)$ Number of exposed individuals in age group $i$
$I_{i}(t)$ Number of infectious individuals in age group $i$
$R_{i}(t)$ Number of recovered individuals in age group $i$
Parameter Value Description
$\beta_{0}$ $3.99\times 10^{-8}$ Transmission rate coefficient (constant across all age groups)
$\sigma$ 0.5 Rate of latent individuals becoming infectious (constant across all age groups)
$\gamma$ 0.5 Recovery rate (constant across all age groups)
Table 3: State variables and parameters of the model by Pitman et al. [37] and utilizing the UK population for simulations of Eq A.1.
Appendix B The Governing Equations of the Covid-19 Model by Röst et al.
In this COVID-19 model (see Fig. 5), individuals begin in the susceptible compartment ($S$), representing those who have not yet been infected but are at risk of exposure. Upon contact with an infectious individual, susceptible individuals transition into the latent phase, progressing through two sequential compartments ($L_{1}$ and $L_{2}$). During this period, they are infected but not yet capable of transmitting the disease. After the latent phase, individuals enter the presymptomatic stage ($I_{p}$), where they become infectious but have not yet developed symptoms.
Figure 5: Diagram of the transmission model from [33] utilized to illustrate the proposed framework. Solid black arrows represent the movement of patients between compartments, whereas dashed arrows show potential pathways for the spread of infection.
From the presymptomatic phase, individuals diverge into two possible infection pathways. A fraction of them remain asymptomatic, progressing through three stages ($I_{a,1}\to I_{a,2}\to I_{a,3}$) before recovering ($R$). Despite not showing symptoms, asymptomatic individuals contribute to transmission. The remaining infected individuals develop symptoms and transition through three symptomatic stages ($I_{s,1}\to I_{s,2}\to I_{s,3}$).
A proportion of symptomatic individuals experience severe illness and require hospitalization, moving into the hospitalized compartment ($I_{h}$). From this state, patients may either recover ($R$) or deteriorate further, requiring intensive care ($I_{c}$). A fraction of critically ill patients progress to a critical state ($I_{cr}$), from which they may either recover ($R$) or die from the disease ($D$).
Notations here are aligned with the parameter file located in the repository of the framework. Here $\mathrm{inf}_{a}$ denotes the relative infectiousness of $I_{a}$ compared to $I_{s}$, for more details about the other parameters and the methodology for parametrization, see [33] and the GitHub repository of this project https://github.com/zsvizi/sensitivity-contact-epidemics/blob/master/data/rost_model_parameters.json. The governing equations for age group $i\in\{0,\dots,15\}$ of the Covid model [33] introduced in Sec. 3.2 take the following form.
${S^{i}}^{\prime}(t)={}$ $-\beta_{0}\frac{S^{i}(t)}{N_{i}}\cdot\sigma_{i}\sum_{k=1}^{16}c_{k,i}\left[I_{p}^{k}(t)+\mathrm{inf}_{a}\sum_{j=1}^{3}I_{a,j}^{k}(t)+\sum_{j=1}^{3}I_{s,j}^{k}(t)\right]$
${L_{1}^{i}}^{\prime}(t)={}$ $\beta_{0}\frac{S^{i}(t)}{N_{i}}\cdot\sigma_{i}\sum_{k=1}^{16}c_{k,i}\left[I_{p}^{k}(t)+\mathrm{inf}_{a}\sum_{j=1}^{3}I_{a,j}^{k}(t)+\sum_{j=1}^{3}I_{s,j}^{k}(t)\right]-2\alpha_{l}L^{i}_{1}(t)$
${L_{2}^{i}}^{\prime}(t)={}$ $2\alpha_{l}L_{1}^{i}(t)-2\alpha_{l}L_{2}^{i}(t),$
${I_{a}^{i}}^{\prime}(t)={}$ $2\alpha_{l}L_{2}^{i}(t)-\alpha_{p}I_{p}^{i}(t)$
${I_{a,1}^{i}}^{\prime}(t)={}$ $p^{i}\alpha_{p}{I}_{p}^{i}(t)-3\gamma_{a}I_{a,1}^{i}(t)$
${I_{a,2}^{i}}^{\prime}(t)={}$ $3\gamma_{a}I_{a,1}^{i}(t)-3\gamma_{a}I_{a,2}^{i}(t)$
${I_{a,3}^{i}}^{\prime}(t)={}$ $3\gamma_{a}I_{a,2}^{i}(t)-3\gamma_{a}I_{a,3}^{i}(t)$ (B.1)
${I_{s,1}^{i}}^{\prime}(t)={}$ $(1-p^{i})\alpha_{p}I_{p}^{i}-3\gamma_{s}I_{s,1}^{i}(t)$
${I_{s,2}^{i}}^{\prime}(t)={}$ $3\gamma_{s}I_{s,1}^{i}(t)-3\gamma_{s}I_{s,2}^{i}(t)$
${I_{s,3}^{i}}^{\prime}(t)={}$ $3\gamma_{s}I_{s,2}^{i}(t)-3\gamma_{s}I_{s,3}^{i}(t)$
${I_{h}^{i}}^{\prime}(t)={}$ $h^{i}(1-\xi^{i})3\gamma_{s}I_{s,3}^{i}(t)-\gamma_{h}I_{h}^{i}(t)$
${I_{c}^{i}}^{\prime}(t)={}$ $h^{i}\xi^{i}3\gamma_{s}I_{s,3}^{i}(t)-\gamma_{c}I_{c}^{i}(t)$
${I_{\mathrm{cr}}^{i}}^{\prime}(t)={}$ $(1-\mu^{i})\gamma_{c}I_{c}^{i}(t)-\gamma_{\mathrm{cr}}I_{\mathrm{cr}}^{i}(t)$
${R^{i}}^{\prime}(t)={}$ $3\gamma_{a}I_{a,3}^{i}(t)+(1-h^{i})3\gamma_{s}I_{s,3}^{i}(t)+\gamma_{h}I_{h}^{i}(t)+\gamma_{\mathrm{cr}}I_{\mathrm{cr}}^{i}(t)$
${D^{i}}^{\prime}(t)={}$ $\mu^{i}\gamma_{c}I_{c}^{i}(t).$
Appendix C Rearranging the contact matrix elements
We establish a link between the pair $(i,j)$ and $k$, and then recover $(i,j)$ from $k$ using specific association rules. These rules are defined as follows:
  • 1.
    If $i=1$, then $(i,j)\mapsto j$.
  • 2.
    If $i\geq 2$, then $(i,j)\mapsto j+\frac{32+(2-i)}{2}\cdot(i-1)$.
To map $k$ to $(i,j)$:
  • 1.
    If $k\leq 16$, then $k\mapsto(1,j)$.
  • 2.
    If $k>16$, identify $i$ such that
    $\frac{32+(2-i)}{2}\cdot(i-1)<k<\frac{32+(2-(i+1))}{2}\cdot i.$ (C.1)
Next, determine $j$ as
$j=k-\frac{32+(2-i)}{2}\cdot(i-1)$ (C.2)
for $1\leq k\leq 136$. This method preserves the symmetry of the total contact matrix derived from the adjusted matrix.